Many employers provide 401K matching That is, once an employee has been with the firm for a sufficient period (often one year), the firm will match the employee’s contributions to the firm’s available 401K plan up to a certain percentage. Assume that an employee has just become eligible for the firm’s matching plan. In this case, the firm will match the employee’s investments up to 6% of the employee’s salary. (If the employee contributes 1%, the company puts in 1%, likewise up through 6%. If the employee puts in more than 6%, the company’s contribution remains at 6% of the salary.) If the 401K investment plan currently has earned an average of 5% per year for the past 5 years, consider the following:

What percentage of their salary should this employee invest in the 401K if they believe that they can earn 8% per year in an ETF (exchange-traded fund)? How does the current investment environment affect this employee’s decision? What might be circumstances you can envision that might change the employee’s investment decisions? Use the time value of money concepts that you have learned in this class to help you consider this problem. Be sure to consider the benefits and risks of investments in general and those for this example in particular. If you decide to use numeric examples, be sure to state your assumptions.

Note: 401K and similar plans allow employees to invest pre-tax funds as savings for their retirement. Taxes are paid when the funds are withdrawn during retirement, with steep penalties for early withdrawal of funds. “Regular” investments are made with after-tax funds.
Answer #1

1. What percentage of their salary should this employee invest in the 401K if they believe that they can earn 8% per year in an ETF (exchange-traded fund)?

Assume employee puts x% of salary in 401K, where x is less than 6%. Company also puts x% on behalf of the employee.

Total returns after 1 year assuming 5% interest rate = 2x% of Salary * 0.05 = 0.02xS * 0.05 = 0.001xS (max return with x=6 is 0.006S)

If x% is put into an ETF generating 8% interest, then total returns after 1 year = x%S * 0.08 = 0.0008xS (max return with x=6 is 0.0048S)

Assume employee puts (6+y)% of salary in 401K. Company puts 6% on behalf of the employee.

Total returns after 1 year assuming 5% interest rate = 0.01 * (12+y) *S * 0.05 = 0.006S+ 0.0005yS

If (6+y)% of salary is put into an ETF generating 8% interest, then total returns after 1 year = 0.01 * (6+y) *S * 0.08 = 0.0048S + 0.0008yS

For indifference, 0.006S+ 0.0005yS = 0.0048S + 0.0008yS

0.0012S = 0.0003yS

y=4

Therefore, total returns after 1 year with y as 4% = (0.006 + 0.002)S = 0.008S

Since 0.008S is greater than the max return possible with less than 6% investment in either 401K or ETF, hence the employee should put total 10% of salary into 401K

2. How does the current investment environment affect this employee’s decision?

Higher rate of ETF provides opportunities to the employee to generate returns higher than a plan vanilla 401K with some amount of risk taking (as ETFs are as volatile as the index)

3. What might be circumstances you can envision that might change the employee’s investment decisions?

If ETFs start generating lower returns than 401K, then the employee will invest all money into 401K and provide no allocation to ETFs